Please can you help me with these and explain step by step thank you:
1) x^2+8+7=0
2)15x^3-51x^2+18x
3) 6x^2+17x+12
4) 147x^2y-27

Answer by mananth(16949)

6x^2+17x+12
compare this equation with ax^2+bx+c
All problems have to be worked out the following way.
coefficient first term =a= 6
coefficient II term = b=17
coefficient III term = c=12
Multiply a&c we get 72
factorize 72 = 2*2*2*3*3
split the factors in two groups such that when you add up the groups you should get 17.
In this case 2*2*2=8
3*3=9
9+8 =17
Rewrite the polynomial
6x^2+9x+8x+12
3x(x+3)+4(x+3)
(x+3)(3x+4)

15x^3-51x^2+18x
facor out the common term
3x(5x^2-17x+6)
Follow the same steps as above
a=5,b=-17,c=6
ac=30, factors = 2*3*5
groups -15 & -2 = -17 which is b
3x(5x^2-15x-2x+6)
3x(5x(x-3)-2(x-3))
3x(x-3)(5x-2)